Circular No. 7136

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE A. Kim, Collège de France, and N. Regnault, Laboratoire de l’Accélér-

ateur Linéaire, Université de Paris-Sud, on behalf of the EROS collaboration; and P. Nugent, Lawrence Berkeley National Laboratory, on behalf of the Supernova Cosmology Project, report the discovery of nine supernovae in the course of the automated supernova search with the 1-m Marly telescope (+ CCD) at La Silla (limiting mag ∼ 21.5; cf. IAUC 7117):

SN  1999 UT  α_{2000}  δ_{2000}  V  Offset
1999bi Mar. 11 11^{h}01^{m}15.^{s}76 −11^{o}45′15′′.2 20.5 0′′.8 E, 0′′.8 N
1999bj Mar. 11 11 51 38.39 −12 29 08.3 20.8 0′′.8 E, 4′′.0 S
1999bk Mar. 15 11 28 52.01 −12 18 08.3 19.1 0′′.2 E, 6′′.5 S
1999bl Mar. 15 11 12 13.60 − 5 04 44.8 21.2 2′′.2 E, 2′′.7 S
1999bm Mar. 15 12 45 00.84 − 6 27 30.2 20.5 3′′.4 E, 1′′.4 N
1999bn Mar. 17 11 57 00.40 −11 26 38.4 19.6 0′′.9 W, 0′′.2 N
1999bo Mar. 18 14 21 07.48 − 5 57 25.8 19.7 7′′.6 E, 2′′.7 N
1999bp Mar. 20 11 39 46.42 − 8 51 34.8 18.8 2′′.6 E, 2′′.6 N
1999bq Mar. 20 13 06 54.46 −12 37 11.6 20.1 1′′.7 E, 4′′.1 N

CCD spectra show that all of these are consistent with type-Ia events; the following data indicate the date the spectrum was obtained, the observer (A = G. Aldering at the Cerro Tololo 4-m telescope; D = T. Dahlen at the Nordic Optical Telescope, with analysis contributed by A. Goobar and I. Hook), the redshift (z), and the estimated distance in time from maximum light: SN 1999bi, Mar. 16, A, 0.124, near; SN 1999bj, Mar. 16, A, 0.16, 17 days after; SN 1999bk, Mar. 16, A, 0.096, 2 days after; SN 1999bl, Mar. 24, A, 0.3, near; SN 1999bm, Mar. 21, A, 0.15, near; SN 1999bn, Mar. 23, D, 0.12, near; SN 1999bo, Mar. 30, A, 0.13, after; SN 1999bp, Mar. 21, A, 0.077, 8 days before; SN 1999bq, Mar. 23, D, 0.149, near.

COMET 37P/FORBES Visual m_{1} estimates: Mar. 21.50 UT, 13.3: (A. Hale, Cloudcroft, NM,

0.41-m refl.); 27.86, 13.5 (A. Pearce, Nedlands, W. Australia, 0.41-m refl.).

COMET 29P/SCHWASSMANN-WACHMANN 1 Total magnitude estimates, visual unless otherwise noted: Mar. 22.77

UT, 15.3: (A. Nakamura, Kuma, Japan, 0.60-m reflector + CCD); 28.48, 13.0: (Hale); 30.84, 12.8 (Pearce); 31.87, 12.6 (Pearce).

1999 April 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green